a force produces an acceleraion of 1,1/2,1/,1/4..........all in m/s2,when combined separately to n bodies to form single one,then the acceleration of the system will be ,if same force is taken into account

Grade:12

1 Answers

Arun
6 years ago
Dear Aniket
 
M = m1 + m2 + m3 + ......
F/a = F/a1 + F/a2 + .....
1/a = 1/a1 + 1/a2 + 1/a3 +.....
1/a = 1 + 2 + 4 + 8 +....
1/a = 1 (2^n -1)/(2-1)
1/a = 2^n  - 1
a = 1/(2^n  -1)